# INR-TH-2023-024 Anomalous cosmic-ray correlations revisited with a complete full-sky sample of BL Lac type objects

### 12 Dec 2023 v1

@@ Abstract
Cosmic rays with energies above $10^{19}$ eV, observed in 1999–2004 by the High Resolution Fly’s Eye (HiRes) experiment in the stereoscopic mode, were found to correlate with directions to distant BL Lac type objects (BL Lacs), suggesting non-standard neutral particles travelling for cosmological distances without attenuation. This effect could not be tested by newer experiments because of their inferior angular resolution. The distribution in the sky of BL Lacs associated with cosmic rays was found to deviate from isotropy, which might give a clue to the interpretation of the observed anomaly. However, previous studies made use of a sample of BL Lacs which was anisotropic by itself, thus complicating these interpretations. Here, we use a recently compiled isotropic sample of BL Lacs and the same HiRes data to confirm the presence of correlations and to strengthen the case for the local large-scale structure pattern in the distribution of the correlated events in the sky. Further tests of the anomaly await new precise cosmic-ray data.
@@
Confirmed in Ref. [3], where unpublished HiRes data were used, the correlations [1] remain neither tested with other experiments nor unambiguously explained. The reason for the lack of tests is in the fact that none of modern experiments reaches the angular resolution of HiRes stereo, where, according to the Monte-Carlo simulations, 68% of events were reconstructed within $0.55^{\circ}$ from their thrown direction, and 95% – within $1.26^{\circ}$ [4]. Newer experiments increased the distance between detectors in order to cover larger effective area, hence worsened their angular resolution. An attempt to test the results of [1, 3] was reported at a conference [5] but was not conclusive, see e.g. Refs. [6, 7] for discussion.
@@
One complication of searches for the anisotropy of correlated events is the intrinsic anisotropy of the BL Lac catalog [13] used in previous studies. Being a compilation of all data available for that moment, it covers different areas of the sky with different and not documented sensitivities. Besides the terrestrial reasons for anisotropy (different fields of view of telescopes), there is a significant Galactic anisotropy caused by the absorption near the Galactic plane (objects in Ref. [1] were selected by their observed visual magnitude, not corrected for the absorption) and by consequent complications of spectral observations required to determine the BL Lac nature of the source. In general, these anisotropies have nothing to do with the LSS and can hardly affect the conclusions of the study, but biases could not be excluded. For instance, one of the suspects is the occasional almost perpendicular orientation of the Galactic and Supergalactic planes.
With the purpose to avoid such hard-to-control biases, we have recently compiled a new catalog of optically selected blazars [14] with the isotropic full-sky coverage. As a part of the project, a sample of confirmed BL Lacs was determined following the same criteria which were used in [13], see Ref. [15] for details. Here, we use this isotropic sample to test the LSS-related pattern [7] in the distribution of BL Lacs correlated with HiRes cosmic rays. We also make some comments regarding future tests of the correlations with new cosmic-ray data.
@@
The sample of BL Lacs we use is described in Ref. [14]. It was constructed on the basis of full-sky catalogs of blazars selected by very-long-baseline radio interferometry and by gamma-ray observations. The selection of optically bright sources, similar to [1], was performed on the basis of the GAIA DR3 $G$-band magnitude, corrected for the Galactic absorption, $G_{\rm corr}<18^{\rm m}$. There are 336 sources in the sample, and their distribution in the sky satisfies the quantitative criteria of isotropy described in Ref. [14].
@@
3.3. Distribution of correlated directions in the sky. We turn now to the main topic of the present study and use the isotropic sample of BL Lacs to search for the LSS pattern in the distribution of correlated cosmic-ray arrival directions in the sky. Following Ref. [7], we use the weighted density of galaxies $f$ from the three-dimensional catalog of the Two Micron All Sky Survey [17, 18] as the LSS template. Details of the construction of the template, performed along the lines of Ref. [19, 20], are described in Ref. [7]. Figure 2
@@
presents the sky map with the correlating BL Lacs shown on top of this template distribution. One can see that, like in Ref. [8], correlating sources tend to fall in the darker shadow regions, corresponding to larger values of $f$. This is confirmed by a statistical study: making use of the Kolmogorov-Smirnov distances between sets of values of $f$ for correlated BL Lacs and for all HiRes events, we obtain the p-value of $3.9\times 10^{-3}$ for 19 BL Lac – cosmic ray pairs to follow the same distribution of $f$ as the entire HiRes data set. The same quantity for the sample used in [1] is $2.6\times 10^{-3}$. Therefore, despite the addition of many new sources to the catalog, the LSS pattern in the distribution of BL-Lac associated events remains the same. Note that this is true only for cosmic rays associated with BL Lacs: for the entire HiRes sample, correlations with LSS were not found [21].
@@
While the low angular resolution of contemporary cosmic-ray observatories prevented one from testing the effects described here with new data, the large statistics collected by these experiments might help to partially overcome this difficulty. In the nearest future, tests of the BL Lac correlations with cosmic rays could be performed, in particular, with the help of the data collected by the Telescope Array experiment [22], which is a successor of HiRes. The present study allows us to make a few comments which might be useful in planning the strategy of these tests.
@@
  • The actual angle $\theta$, at which the correlations are most significant, depends on the fraction of cosmic-ray events coming from BL Lacs and on the fraction of BL Lacs in the sample, which emit cosmic rays. Both numbers are small and fluctuate, so it makes sense to either consider various values of $\theta$.
@@
Acknowledgements. The authors are indebted to M. Kuznetsov and G. Rubtsov for interesting and helpful discussions. This work was supported by the Russian Science Foundation, grant 22-12-00253.
@@
  • [3] Abbasi, R. U., T. Abu-Zayyad, T., Amann, J. F., et al. Astrophys. J. 636, 680 (2006).
@@
  • [14] Kudenko, M. and Troitsky, S. Accompanying paper in today’s arXiv (2023).
@@
  • [21] Abbasi, R. U. et al. Astrophys. J. Lett. 713, L64–L68 (2010).
  • [22] Kawai, H. et al. Nucl. Phys. B Proc. Suppl. 175-176, 221 (2008).

### 23 Jan 2024 v3

@@ Abstract
Cosmic rays with energies above $10^{19}$ eV, observed in 1999–2004 by the High Resolution Fly’s Eye (HiRes) experiment in the stereoscopic mode, were found to correlate with directions to distant BL Lac type objects (BL Lacs), suggesting non-standard neutral particles travelling for cosmological distances without attenuation. This effect could not be tested by newer experiments because of their inferior angular resolution. The distribution in the sky of BL Lacs associated with cosmic rays was found to deviate from isotropy, which might give a clue to the interpretation of the observed anomaly. However, previous studies made use of a sample of BL Lacs which was anisotropic by itself, thus complicating these interpretations. Here, we use a recently compiled isotropic complete sample of BL Lacs and the same HiRes data to confirm the presence of correlations and to strengthen the case for the local large-scale structure pattern in the distribution of the correlated events in the sky. Further tests of the anomaly await new precise cosmic-ray data.
@@
Confirmed in Ref. [3], where unpublished HiRes data were used, with the same pre-trial p-value of $2\times 10^{-4}$, corresponding to the post-trial $p\sim 10^{-3}$, the correlations [1] remain neither tested with other experiments nor unambiguously explained. The reason for the lack of tests is in the fact that none of modern experiments reaches the angular resolution of HiRes stereo, where, according to the Monte-Carlo simulations, 68% of events were reconstructed within $0.55^{\circ}$ from their thrown direction, and 95% – within $1.26^{\circ}$ [4]. Newer experiments increased the distance between detectors in order to cover larger effective area, hence worsened their angular resolution. The only attempt to test the results of [1, 3] was reported at a conference [5] but was not conclusive. Indeed, Ref. [6] estimated the numbers of cosmic-ray events required to test the correlations [1] by different experiments. For the surface detector of the Pierre Auger Observatory, this number was about 3500, while only 1672 events were used in Ref. [5]. See e.g. Refs. [7, 8] for discussion of other complications related to these tests.
@@
One complication of searches for the anisotropy of correlated events is the intrinsic anisotropy of the incomplete BL Lac catalog [14] used in previous studies. Being a compilation of all data available for that moment, it covers different areas of the sky with different and not documented sensitivities. Besides the terrestrial reasons for anisotropy (different fields of view of telescopes), there is a significant Galactic anisotropy caused by the absorption near the Galactic plane (objects in Ref. [1] were selected by their observed visual magnitude, not corrected for the absorption) and by consequent complications of spectral observations required to determine the BL Lac nature of the source. In general, these anisotropies have nothing to do with the LSS and can hardly affect the conclusions of the study, but biases could not be excluded. For instance, one of the suspects is the occasional almost perpendicular orientation of the Galactic and Supergalactic planes: because of this, the fraction of directions falling in the zone of avoidance is smaller for the local filament, compared to the full HiRes field of view.
With the purpose of avoiding such hard-to-control biases, we have recently compiled a new catalog of optically selected blazars [15] with the isotropic full-sky coverage. As a part of the project, a sample of confirmed BL Lacs was determined following the same criteria which were used in [14], see Ref. [16] for details. Here, we use this isotropic sample to test the LSS-related pattern [8] in the distribution of BL Lacs correlated with HiRes cosmic rays. We also make some comments regarding future tests of the correlations with new cosmic-ray data.
@@
The sample of BL Lacs we use is described in Ref. [15]. It was constructed on the basis of full-sky catalogs of blazars selected by very-long-baseline radio interferometry (VLBI) and by gamma-ray observations. The sample is a combination of two complete samples, flux-limited at 8 GHz VLBI and above 1 GeV, respectively.
The selection of optically bright sources, motivated by the cuts used in Ref. [1] and not related to the completeness of the isotropic sample, was performed on the basis of the GAIA DR3 $G$-band magnitude, corrected for the Galactic absorption, $G_{\rm corr}<18^{\rm m}$. There are 336 sources in the sample, and their distribution in the sky satisfies the quantitative criteria of isotropy described in Ref. [15].
@@
3.3. Distribution of correlated directions in the sky. We turn now to the main topic of the present study and use the isotropic sample of BL Lacs to search for the LSS pattern in the distribution of correlated cosmic-ray arrival directions in the sky. Following Ref. [8], we use the weighted density of galaxies $f$ from the three-dimensional catalog of the Two Micron All Sky Survey [18, 19] as the LSS template. Details of the construction of the template, performed along the lines of Ref. [20, 21], are described in Ref. [8]. Note that all BL Lacs are located far beyond the LSS, at distances between $\sim 150$ Mpc and a few gigaparsecs, while typical distances contributing to the template are $\lesssim 30$ Mpc. Correlations with sources located that far justifies the neglection of any attenuation at $\sim 30$ Mpc assumed in the calculation of $f$ in Ref. [8].
Figure 2
@@
presents the sky map with the correlating BL Lacs shown on top of this template distribution. One can see that, like in Ref. [9], correlating sources tend to fall in the darker shadow regions, corresponding to larger values of $f$. This is confirmed by a statistical study: making use of the Kolmogorov-Smirnov distances between sets of values of $f$ for correlated BL Lacs and for all HiRes events, we obtain the p-value of $3.9\times 10^{-3}$ for 19 BL Lac – cosmic ray pairs, closer than the 95% CL angular resolution, to follow the same distribution of $f$ as the entire HiRes data set. The same quantity for the sample used in [1] is $2.6\times 10^{-3}$. Therefore, despite the addition of many new sources to the catalog, the LSS pattern in the distribution of BL-Lac associated events remains the same. Note that this is true only for cosmic rays associated with BL Lacs: for the entire HiRes sample, correlations with LSS were not found [22].
@@
While the low angular resolution of contemporary cosmic-ray observatories prevented one from testing the effects described here with new data, the large statistics collected by these experiments might help to partially overcome this difficulty. In the nearest future, tests of the BL Lac correlations with cosmic rays could be performed, in particular, with the help of the data collected by the Telescope Array experiment [23], which is a successor of HiRes. To precisely evaluate the sensitivity of the expected analysis and refine the rough estimates of Ref. [6], dedicated simulations incorporating the experiment’s specific details are essential.The present study allows us to make a few comments which might be useful in planning the strategy of these tests.
@@
  • The actual angle $\theta$, at which the correlations are most significant, depends on the fraction of cosmic-ray events coming from BL Lacs and on the fraction of BL Lacs in the sample, which emit cosmic rays. Both numbers are small and fluctuate, so it makes sense to consider various values of $\theta$.
@@
Acknowledgements. The authors are indebted to M. Kuznetsov and G. Rubtsov for interesting and helpful discussions, and to three anonymous referees for useful comments. This work was supported by the Russian Science Foundation, grant 22-12-00253.
@@
  • [3] Abbasi, R. U., Abu-Zayyad, T., Amann, J. F., et al. Astrophys. J. 636, 680 (2006).
@@
  • [15] Kudenko, M. and Troitsky, S. arXiv:2312.07508 (2023).
@@
  • [22] Abbasi, R. U., Abu-Zayyad, T., Allen, M., et al. Astrophys. J. Lett. 713, L64–L68 (2010).
  • [23] Kawai, H., Yoshida, S., Yoshii, H., et al. Nucl. Phys. B Proc. Suppl. 175-176, 221 (2008).